# Architecture Decoupling Is Not All You Need For Unified Multimodal Model

### v2

@@ Abstract
00footnotetext: The work was conducted during the internship of Dian Zheng at Meituan. Project Leader. 🖂corresponding authors.  Code is available
Unified multimodal models for image generation and understanding represent a significant step toward AGI and have attracted widespread attention from researchers. The main challenge of this task lies in the difficulty in establishing an optimal training paradigm due to inherent conflicting targets in understanding and generation tasks. To alleviate these conflicts and pursue higher performance, many researchers adopt varying degrees of model decoupling (e.g., Double image encoders, MOE/MOT architecture, or frozen MLLM). However, excessive model decoupling can lead to the loss of interleave generation ability, undermining the original intent of unified models. In this work, we aim to explore how to mitigate task conflicts without resorting to model decoupling. Firstly, we analyze why decoupling alleviates conflicts by studying the cross-modal attention behavior of models. We observe that model decoupling essentially drives models toward task-specific multimodal interaction patterns, as seen in Qwen-VL and HunyuanImage, and that the more thorough the decoupling, the more consistent the behavior becomes. Motivated by this observation, we propose Attention Interaction Alignment (AIA) loss, which explicitly learns Task-Specific multimodal interaction patterns during training. To demonstrate the generalizability of our AIA loss, we apply it to Emu3 and Janus-Pro during SFT and post-training stage respectively. Without bells and whistles, AIA not only refines cross-modal attention patterns, but also boosts both generation and understanding performance.
@@ 1 Introduction
Although the original intention behind UMM is admirable, practical realities are harsh: visual understanding and generation tasks require distinct feature granularities and representations at different network layers. Early research [39, 43, 33] attempted fully unified architectures (e.g., sharing the image encoder and base model), but the results lagged significantly behind single-task approaches. To address task conflicts and boost performance, some [5, 42, 20, 7] have begun to decouple model components to varying degrees, achieving promising results. Due to the inherent effectiveness of this strategy, more researchers are now pursuing decoupled architectures. However, this trend overlooks the core motivation of UMM [54]: leveraging the unified model’s capacity for cross-modal reasoning to enhance single-task performance. Excessive decoupling risks losing this synergistic benefit, limiting the model’s capability to transfer knowledge and generalize across tasks. Furthermore, architecture decoupling, such as using double image encoders, forces the cross-modal reasoning process to undergo additional decode-encode steps, which is inelegant, time-costing and can result in information loss.
@@
To maintain the original intent of UMM while narrowing its performance gap with decoupled models, we first conducted a detailed analysis of the underlying causes of this gap. Since the core of interaction between generation and understanding tasks lies in cross-modal information exchange, we focused our investigation on cross-modal interaction patterns (see Sec 3.1 for details). As illustrated in Fig 1, we first observe that regardless of architecture decoupling degrees, the two tasks show consistent negative correlation in cross-modal interaction patterns within each layer. We further verified that this phenomenon is independent of input type or length; rather, the model dynamically allocates cross-modal representational weights within layers based on task requirements. Moreover, as decoupling increases, the interaction patterns increasingly resemble those of single-task models. This suggests that existing decoupled models do not eliminate the inherent conflict between tasks; instead, they make each task behave more like its single-task counterpart, resulting in performance improvement.
@@
Based on this observation, we propose Attention Interaction Alignment (AIA) loss, which will explicitly constrain the layer-wise cross-modal interaction intensity during training without architecture decoupling. Specifically, we first select models suitable as cross-modal interaction learning target. The chosen models must demonstrate both high performance and autoregressive capability to meet the requirements of UMM. For the understanding task, we use Qwen3-VL-8B. For generation, since SOTA methods rely on external diffusion heads and purely autoregressive approaches are less effective, we refer to HunyuanImage-3.0 [2], which combines an autoregressive backbone with a diffusion head. Although its generative ability primarily depends on the diffusion head, its attention patterns remain informative. For both understanding and generation tasks, we use 100 samples to extract the attention patterns at each layer and compute their average. Recognizing that attention patterns are closely linked to model architecture and pretraining, we apply the Huber loss to relax layer-wise attention constraints, enabling more flexible allocation of attention weights.
@@
To validate the effectiveness of our proposed AIA loss, we adopt it to two degrees of decoupling methods (e.g., Emu3 for purely unified model, Janus-Pro for slight decoupling model). Experimental results demonstrate that our approach enhances both generation and understanding performance while narrowing the gap with more strongly decoupled methods. Although we can not surpass the performance of highly decoupled methods yet, we believe that it is a step to the real UMMs and our method will be fully open-sourced to help the development of the community. We highlight the main contributions of this paper below:
@@
  • We provide the first mechanistic analysis of unified models with different decoupling architectures through cross-modal attention interaction intensity, revealing that decoupling does not resolve task conflicts but merely shifts attention patterns toward single-task behaviors.
@@ 2 Related Works
2.1 Uunified Multimodal Models (UMMs)
In recent years, with the development of LLM and MLLM [1, 8, 19, 34, 48, 23, 24], autoregressive architectures have been thoroughly explored, demonstrating exceptional capabilities in large models. Researchers are now considering whether MLLMs can be integrated with image generation to form a unified model, enabling automatic interleaved reasoning at the latent level. Initial unified MLLMs, such as Liquid [43], Emu3 [39], and Chameleon [33], adhered strictly to this path but used VAE [17, 9, 36] as the image encoder, limiting their understanding capabilities. To address tokenizer representation conflicts, methods like UniLip [32] and TokenFlow [29] have constructed a unified tokenizer compatible with both understanding and generation (coarse-grained and fine-grained) feature requirements. The Janus series [5, 40, 26] and show-o [46, 47] series have attempted to decouple the image encoder and training objectives for generation and understanding, achieving further performance improvements. However, due to conflicts between understanding and generation in the backbone network, performance remains constrained. To alleviate task conflicts, approaches like BAGEL [7] and OneCat [20] have explored partially decoupled architectures (MOT, MOE), yielding promising results. Meanwhile, another group of researchers argues that generation hardly aids understanding, so Metaquery [27], OmniGen2 [42], UniWorld-V1 [22], and Blip3-o [4] have chosen to fix the MLLM and solely optimize the diffusion head, achieving current state-of-the-art results. However, as model architectures become increasingly decoupled, starting from the image encoder, models can no longer achieve automatic interleaved reasoning in the latent space but must undergo an inelegant decode-encode process. Furthermore, as the backbone network becomes decoupled, the model’s ability to handle understanding and generation in a unified manner is further weakened, let alone completely fixing the MLLM, which has deviated from the original intent of UMM. This paper aims to explore why model decoupling can alleviate conflicts and enhance performance while attempting to achieve similar effects within a fully unified architecture.
@@ 2.2 Ultra Task-Specific Models
Through extensive technological iterations, understanding tasks have developed into a nearly fixed model architecture. Models like Qwen [1, 34, 48] and the LLaVA [8, 19] series employ a semantic encoder combined with an autoregressive architecture, achieving outstanding results. In contrast, generation tasks lack a fixed model architecture. Early models such as SDXL [28], SD3 [10], and the FLUX [18] series utilized CLIP [30] as the text encoder within a pure diffusion [15, 31] framework, achieving high aesthetic quality but showing some limitations in instruction compliance. SimpleAR [38] explored image generation using LLM as the base model within a purely autoregressive architecture, but the inherent information loss in discrete representations resulted in images with a pronounced blur. Qwen-Image [41], and HunyuanImage-3.0 [2] combined the strengths of both approaches by replacing CLIP with MLLM and integrating a diffusion head after the MLLM, achieving excellent results in both instruction compliance and aesthetic quality. In this paper, we select the most performant models in understanding and generation tasks as references for cross-modal interaction patterns, thereby enhancing the reliability of the conclusions.
@@ 3.1 Analysis of Cross-Modal Attention Interaction Pattern in Various Model Architectures
Method. We take text-to-image generation as an example. As shown in Fig 2, after obtaining the attention map from the softmax operation, we calculate the sum of all text tokens in each row of the attention map. Then, we compute the average over all layers, all image tokens and all attention heads. This value represents the cross-modal interaction intensity between images and text for the given sample. The specific formula is as follows:
@@
where I means the intensity, l is the layer index, H, Q, K represent head, query, key numbers respectively. The average results of 100 samples are shown in Fig 1. In Table 1, we further validate that the attention pattern of each model is independent of input length and type (i.e., consisting of prompts like text rendering, short and dense caption and questions like caption, choices, etc.) by computing the standard deviation for each model.
@@
Rationality of task-specific attention pattern. Fig 1 shows that Qwen3-VL consistently exhibits low attention to image tokens, which aligns with the motivation behind token pruning methods [37, 51] in current understanding tasks. For generation, HunyuanImage maintains around 40% attention to text tokens in the first 80 layers, with a sharp decline in the final layers. This pattern matches the common consensus in generative models: shallow layers focus on building semantic representations and thus attend more to high-level text features, while the final layers shift toward pixel-level image features, resulting in reduced attention to text. The decrease in text attention in deeper layers further supports the claim in RecA [45] that textual information is insufficient at these stages. Except for Emu3 [39], all other architectures follow the single-task trend, indicating that unified models tend to learn in a task-specific manner once their architectures are decoupled.
@@
[…]
SDXL [28] - Diff - 0.55 74.65
SD3-medium [10] 2B Diff - 0.74 84.08
Infinity [14] 2B VAR - 0.73 83.50
Infinity [14] 8B VAR - 0.79 86.60
FLUX.1-dev [18] 12B Diff - 0.82 84.00
Emu3-Gen [39] 8B AR - 0.66 80.60
Qwen-Image [41] 7B+20B AR+Diff - 0.87 88.32
[…]
Emu3-Chat [39] 8B AR 31.6 58.5 36.6 37.2 85.2 1244 - -
Qwen2.5-VL [1] 3B AR 53.1 79.1 - 61.8 - - - -
Qwen2.5-VL [1] 7B AR 58.6 83.5 - 66.6 - 1685 - -
InternVL2.5 [6] 8B AR 56.2 84.6 - 62.8 90.6 - - -
InternVL3 [53] 8B AR 62.7 83.4 - 81.3 91.1 - - -
Qwen3-VL 8B AR 69.6 85.0 - - - - - -
[…]
MetaQuery-XL [27] 7B+1.6B AR+Diff 58.6 83.5 - 66.6 - 1685 0.80 82.05
Blip3-o [4] 7B+1.4B AR+Diff 58.6 83.5 - 66.6 - 1685 0.84 81.60
UniWorld-V1 [22] 7B+12B AR+Diff 58.6 83.5 - 67.1 - 1685 0.84 81.38
OmniGen2 [42] 3B+4B AR+Diff 53.1 79.1 - 61.8 - - 0.86 83.57
[…]
BAGEL [7] 7B+7B AR+Diff 55.3 85.0 69.3 67.2 - 1687 0.88 85.07
OneCat [20] 3B+3B+3B AR+VAR 41.9 78.8 - 52.2 - 1630 0.90 84.53
[…]
Show-o [46] 1.3B AR+Mask 26.7 - - - 80.0 1097 0.69 67.27
Show-o2 [47] 7B AR+Diff 48.9 79.3 - - - 1621 0.76 86.14
Janus-Pro [5] 7B AR 41.0 65.54 (79.2) 47.3 50.0 87.4 1567 0.80 84.19
@@
[…]
Chameleon [33] 7B AR 28.4 35.7 - 8.3 - - 0.39 -
VILA-U [44] 7B AR 32.2 66.6 22.0 27.7 83.9 1336 0.39 72.48
Emu3 [39] $\dagger$ 8B AR 31.6 61.4 8.7 15.1 77.3 910 0.60 79.24
@@
Why model decoupling improves performance. From Table 1, we can first rule out the influence of input properties on cross-modal attention patterns. Fig 1 shows that the interaction curves for the two tasks have negative correlation, indicating that, after unified training, the inherent conflict between tasks forces the model to allocate attention weights to different tasks at different layers. This mechanism helps the model self-mitigate cross-task interference, and the resulting allocation pattern becomes fixed. Furthermore, Fig 1 demonstrates that this negative correlation persists regardless of the degree of model decoupling. As the model becomes more decoupled, the attention interaction patterns increasingly resemble those of task-specific models, leading to improved performance. Specifically, when performing understanding tasks, Janus-Pro exhibits relatively high attention to image tokens in the initial layers, which limits its understanding capability. For Emu3, both generation and understanding tasks deviate significantly from the typical single-task interaction patterns, highlighting the inherent difficulty of fully unified learning within an autoregressive architecture.
3.2 Attention Interaction Alignment Loss
Based on the observation above, we propose attention interaction alignment loss, which will constrain the attention patterns explicitly during training. Specifically, we use the layer-wise intensity from the task-specific models in Fig 1 as learning targets, termed $T_{l}$. However, since the attention curves are fixed throughout network training, it is unsuitable to apply overly strict constraints for supervision. Therefore, we divide the values into several sub-stages according to their magnitudes and further relax the absolute constraint on individual values using the Huber loss. Taking Emu3 [39] as an example, the specific formula is as follows:
where $I_{l}$ denotes the cross-modal intensity at layer $l$, and for the layer-wise target boundary $T_{l}$ and Huber threshold $\delta_{l}$, please see the Supplementary File. The AIA loss will combine with next-token-prediction loss, term $\mathcal{L}_{\text{NTP}}$ as:
where $\lambda$ is set to 40 in the experiment by default.
@@ 3.2 Attention Interaction Alignment Loss
w/o stage-level intensity 40.2 67.4 43.3 44.0 87.6 1543.2 0.79 83.79
w/o Huber 41.2 73.1 47.6 47.7 88.2 1613.9 0.80 84.15
w/o AIA (baseline) 40.7 71.5 47.1 49.2 88.1 1593.1 0.80 84.15
@@
Qwen3-VL+FLUX 40.2 72.1 45.7 46.0 87.5 1555.1 0.76 82.89
Qwen3-VL+SimpleAR 41.5 74.3 47.9 48.6 89.5 1623.4 0.80 84.12
@@
Qwen3-VL+Qwen-Image 40.3 72.5 44.3 46.6 88.0 1546.7 0.75 83.01
@@
Gen:Und: 2:1 41.8 74.2 46.9 49.3 89.1 1643.6 0.81 84.23
Gen:Und: 4:1 41.2 72.8 47.5 48.5 88.3 1621.4 0.80 83.87
Gen:Und: 1:2 42.0 74.8 47.7 49.3 89.3 1645.9 0.79 83.12
3.3 Training Details Under Different Architectures
Emu3. We load the pretrained (PT) weights of Emu3 and perform SFT training using our own data. Note that Emu3 only employs unified training at this stage; the results reported in the original paper were obtained through separate training. Since Emu3’s performance at the pretraining stage is relatively poor, as shown in Fig 3, we incorporate the AIA loss with varying weights during the SFT stage and observe that the NTP loss convergence trends remain nearly identical. This indicates that, when the learning target is set appropriately, incorporating the AIA loss during SFT does not significantly affect the model’s pretrained knowledge.
Janus-Pro. Since Janus-Pro only provides the final SFT weights, we perform post-training on this basis. When data quality does not differ significantly, this greatly increases the difficulty of tuning, as the model’s distribution is already highly fixed at this stage (i.e., a point further supported by the variance comparison between Emu3 and Janus-Pro in Table 1). However, this also aligns more closely with realistic settings: how to perform fine-grained tuning when only the final weights are accessible and attention distribution adjustment is required. As shown in Fig 3, the model is highly sensitive to the weight of the AIA loss at this stage, yet it can still achieve the desired effect when $\lambda$ is set appropriately. We further validate this in Section 4.3.

### v3

# AIA: Rethinking Architecture Decoupling Strategy In Unified Multimodal Model

@@ Abstract
Unified multimodal models for image generation and understanding represent a significant step toward AGI and have attracted widespread attention from researchers. The main challenge of this task lies in the difficulty in establishing an optimal training paradigm due to inherent conflicting targets in understanding and generation tasks. To alleviate these conflicts and pursue higher performance, many researchers adopt varying degrees of architecture decoupling (e.g., Double image encoders, MOE/MOT architecture, or frozen MLLM). However, excessive model decoupling can lead to the loss of interleave generation ability, undermining the original intent of unified models. In this work, we aim to explore how to mitigate task conflicts without resorting to model decoupling. Firstly, we analyze why decoupling boosts performance by studying the cross-modal attention behavior of models. We observe that architecture decoupling does not solve task conflicts, but essentially drives models toward cross-modal interaction patterns of task-specific models, as seen in Qwen3-VL and HunyuanImage-3.0, and that the more thorough the decoupling, the more consistent the behavior becomes. Motivated by this observation, we propose Attention Interaction Alignment (AIA) loss, which explicitly learns task-specific multimodal interaction patterns during training. To demonstrate the generalizability of our AIA loss, we apply it to Emu3 and Janus-Pro during SFT and post-training stage respectively. Without bells and whistles, AIA not only refines cross-modal attention patterns, but also boosts both generation and understanding performance.
@@ 1 Introduction
Although the original intention behind UMM is admirable, practical realities are harsh: visual understanding and generation tasks require distinct feature granularities and representations at different network layers. Early research Wang et al. (2024); Wu et al. (2024b); Team (2024) attempted fully unified architectures (e.g., sharing the image encoder and base model), but the results lagged significantly behind single-task approaches. To address task conflicts and boost performance, some researches Chen et al. (2025c); Wu et al. (2025b); Li et al. (2025); Deng et al. (2025) have begun to decouple model components to varying degrees, achieving promising results. Due to the inherent effectiveness of this strategy, more researchers are now pursuing decoupled architectures. However, this trend overlooks the core motivation of UMM Zou et al. (2025); Yang et al. (2025b): leveraging the capacity of the unified model for cross-modal reasoning to enhance single-task performance. Excessive decoupling risks losing this synergistic benefit, limiting the model’s ability to transfer knowledge and generalize across tasks Geng et al. (2025). Furthermore, architecture decoupling, such as using double image encoders, forces the cross-modal reasoning process to undergo additional decode-encode steps, which is inelegant and time-costing.
@@
To maintain the original intent of UMM while narrowing its performance gap with decoupled models, we first conducted a detailed analysis of the underlying causes of this gap. Since the core of interaction between generation and understanding tasks lies in cross-modal information exchange, we focused our investigation on cross-modal interaction patterns (see section˜3.1 for details). As illustrated in fig.˜1, we first observe that regardless of architecture decoupling degrees, the two tasks show consistent negative correlation in cross-modal interaction patterns within each layer. We further verified that this phenomenon is independent of input type or length; rather, the model dynamically allocates cross-modal representational weights within layers based on task requirements. Moreover, as decoupling increases, the interaction patterns increasingly resemble those of task-specific models. This suggests that existing decoupled models do not eliminate the inherent conflict between tasks; instead, they make each task behave more like its single-task counterpart, resulting in performance improvement.
@@
Based on this observation, we propose Attention Interaction Alignment (AIA) loss, which will explicitly constrain the layer-wise cross-modal interaction intensity during training without architecture decoupling. Specifically, we first select models suitable as cross-modal interaction learning target. For the understanding target, we use Qwen3-VL-8B Bai et al. (2025a). For generation target, we refer to HunyuanImage-3.0 Cao et al. (2025), which demonstrates superior generative capabilities despite its unified training paradigm. For both understanding and generation tasks, we use 100 samples to extract the attention patterns at each layer and compute their average. Recognizing that attention patterns are closely linked to model architecture and pretraining, we apply the Huber loss to relax layer-wise attention constraints, enabling more flexible allocation of attention weights.
@@
To validate the effectiveness of our proposed AIA loss, we adopt it to two degrees of decoupling methods (e.g., Emu3 for purely unified model, Janus-Pro for slight decoupling model). Experimental results demonstrate that our approach enhances both generation and understanding performance while narrowing the gap with more strongly decoupled methods. We further confirm that the performance gain stems from the alignment of attention patterns rather than mere knowledge distillation, thereby corroborating our initial motivation. We highlight the main contributions of this paper below:
@@
  • We provide the first mechanistic analysis of unified models with different decoupling architectures through cross-modal attention interaction intensity, revealing that decoupling does not resolve task conflicts but merely shifts attention patterns toward task-specific behaviors.
@@ 2 Related Works
2.1 Unified Multimodal Models (UMMs)
In recent years, with the development of LLM and MLLM Bai et al. (2025b); Dubey et al. (2024); Li et al. (2024); Team and others (2024); Yang et al. (2025a); Liu et al. (2024a, 2025), autoregressive architectures have been thoroughly explored, demonstrating exceptional capabilities in large models. Researchers are now considering whether MLLMs can be integrated with image generation to form a unified model, enabling automatic interleaved reasoning at the latent level. Initial unified MLLMs, such as Liquid Wu et al. (2024b), Emu3 Wang et al. (2024), and Chameleon Team (2024), adhered strictly to this path but used VAE Kingma and Welling (2013); Esser et al. (2021); Van Den Oord et al. (2017) as the image encoder, limiting their understanding capabilities. To address tokenizer representation conflicts, methods like TokenFlow Qu et al. (2025), Unitok Ma et al. (2025), Atoken Lu et al. (2025), UniLip Tang et al. (2025) have constructed a unified tokenizer compatible with both understanding and generation (coarse-grained and fine-grained) feature requirements. VTP Yao et al. (2025) further scales the unified tokenizer to a significant extent. The Janus series Chen et al. (2025c); Wu et al. (2024a); Ma et al. (2024), show-o Xie et al. (2025b, c) series, and Transfusion Zhou et al. (2024); Cao et al. (2025) series have attempted to decouple the image encoder and training objectives for generation and understanding, achieving further performance improvements. However, due to conflicts between understanding and generation in the backbone network, performance remains constrained. To alleviate task conflicts, approaches like BAGEL Deng et al. (2025) and OneCat Li et al. (2025) have explored partially decoupled architectures (MOT, MOE), yielding promising results. Meanwhile, another group of researchers argues that generation hardly aids understanding, so Metaquery Pan et al. (2025), OmniGen2 Wu et al. (2025b), UniWorld-V1 Lin et al. (2025), and Blip3-o Chen et al. (2025a) have chosen to fix the MLLM and solely optimize the diffusion head, achieving current state-of-the-art results.
However, as model architectures become increasingly decoupled, starting from the image encoder, models can no longer achieve automatic interleaved reasoning in the latent space but must undergo an inelegant decode-encode process. Furthermore, as the backbone network becomes decoupled, the model’s ability to handle understanding and generation in a unified manner is further weakened, let alone completely fixing the MLLM, which has deviated from the original intent of UMM. This paper aims to explore why architecture decoupling can alleviate conflicts and enhance performance, and then attempts to achieve similar effects within a fully unified architecture.
@@ 2.2 Ultra Task-Specific Models
Through extensive technological iterations, understanding tasks have developed into a nearly fixed model architecture. Models like Qwen Bai et al. (2025b); Team and others (2024); Yang et al. (2025a); Bai et al. (2025a) and the LLaVA Dubey et al. (2024); Li et al. (2024) series employ a semantic encoder combined with an autoregressive architecture, achieving outstanding results. In contrast, generation tasks lack a fixed model architecture. Early models such as SDXL Podell et al. (2023), SD3 Esser et al. (2024), and the FLUX Labs (2024) series utilized CLIP Radford et al. (2021) as the text encoder within a pure diffusion Ho et al. (2020); Song et al. (2020) framework, achieving high aesthetic quality but showing some limitations in instruction compliance. SimpleAR Wang et al. (2025b) explored image generation using LLM as the base model within a purely autoregressive architecture, but the inherent information loss in discrete representations resulted in images with a pronounced blur. Qwen-Image Wu et al. (2025a), Longcat-Image Team et al. (2025) and Z-Image Cai et al. (2025) combined the strengths of both approaches by replacing CLIP with MLLM and integrating a diffusion head after the MLLM, achieving excellent results in both instruction compliance and aesthetic quality. HunyuanImage-3.0 Cao et al. (2025) is trained in a unified manner while showing great generation ability. In this paper, we select the most performant models in understanding (Qwen3-VL) and generation (HunyuanImage-3.0) tasks as references for cross-modal interaction patterns, thereby enhancing the reliability of the conclusions.
2.3 Knowledge Distillation
Knowledge distillation Hinton et al. (2015), initially proposed by Hinton, is a technique that transfers knowledge from a teacher model to a student model. FitNets Romero et al. (2014) further advances this approach by incorporating intermediate features to better align teacher and student models. Our AIA loss differs fundamentally from traditional output- or feature-level knowledge distillation in two key aspects. 1) We rely solely on the layer-wise cross-modal interaction intensity. This is a pre-computed statistic that involves no specific feature and eliminates the need for a concurrent teacher model during training. 2) Standard distillation inherits the severe conflict between generation and understanding inherent in UMMs. Consequently, imposing rigid constraints proves counterproductive and degrades performance, as we verify in Table˜4. AIA emerges as an intuitive and effective solution derived directly from our empirical insights into these conflicts.
@@ 3.1 Analysis of Cross-Modal Attention Interaction Pattern in Various Model Architectures
Method. We take text-to-image generation as an example. As shown in fig.˜2, after obtaining the attention map from the softmax operation, we calculate the sum of all text tokens in each row of the attention map. Then, we compute the average over all layers, all image tokens and all attention heads. This value represents the cross-modal interaction intensity between images and text for the given sample. The formula is as follows:
@@
where I means the intensity, l is the layer index, H, Q, K represent head, query, key numbers respectively. The average results of 100 samples are shown in fig.˜1. In Table˜1, we further validate that the attention pattern of each model is independent of input length and type (i.e., consisting of prompts like text rendering, short and dense caption and questions like caption, choices, etc.) by computing the standard deviation for each model.
@@
Rationality of task-specific attention pattern. fig.˜1 shows that Qwen3-VL Bai et al. (2025a) consistently exhibits low attention to image tokens, which aligns with the motivation behind token pruning methods Wang et al. (2025a); Zhang et al. (2024) in current understanding tasks. For generation, HunyuanImage-3.0 Cao et al. (2025) maintains around 40% attention to text tokens in the first 80 layers, with a sharp decline in the final layers. This pattern matches the common consensus in generative models Yu et al. (2025): shallow layers focus on building semantic representations and thus attend more to high-level text features, while the final layers shift toward pixel-level image features, resulting in reduced attention to text. The decrease in text attention in deeper layers further supports the claim in RecA Xie et al. (2025a) that textual information is insufficient at these stages. Except for Emu3 Wang et al. (2024), all other architectures follow the single-task trend, indicating that unified models tend to learn in a task-specific manner once their architectures are decoupled.
Why model decoupling improves performance. From Table˜1, we can first rule out the influence of input properties on cross-modal attention patterns. fig.˜1 shows that the interaction curves for the two tasks have negative correlation, indicating that, after unified training, the inherent conflict between tasks forces the model to allocate attention weights to different tasks at different layers. This mechanism helps the model self-mitigate cross-task interference, and the resulting allocation pattern becomes fixed. Furthermore, fig.˜1 demonstrates that this negative correlation persists regardless of the degree of model decoupling. As the model becomes more decoupled, the attention interaction patterns increasingly resemble those of task-specific models, leading to improved performance. Specifically, when performing understanding tasks, Janus-Pro exhibits relatively high attention to image tokens in the initial layers, which limits its understanding capability. For Emu3, both generation and understanding tasks deviate significantly from the typical single-task interaction patterns, highlighting the inherent difficulty of fully unified learning within an autoregressive architecture.
3.2 Attention Interaction Alignment Loss
Based on the observation above, we propose attention interaction alignment loss, which will constrain the attention patterns explicitly during training. Specifically, we use the layer-wise intensity from the task-specific models in fig.˜1 as learning targets, termed $T_{l}$. However, since the attention curves are fixed throughout network training, it is unsuitable to apply overly strict constraints for supervision. Therefore, we divide the values into several sub-stages according to their magnitudes and further relax the absolute constraint on individual values using the Huber loss. Taking Emu3 Wang et al. (2024) as an example, the specific formula is as follows:
where $I_{l}$ denotes the cross-modal intensity at layer $l$, and for the layer-wise
Table 2: The hyper-parameter of Huber loss in Emu3 and Janus-Pro.
$0\leq l<10$ $(0.2,0.4)$ $(0.05,0.1)$
$10\leq l<20$ $(0.1,0.4)$ $(0.05,0.15)$
$20\leq l<25$ $(0.1,0.4)$ $(0.05,0.3)$
$25\leq l\leq L$ $(0.05,0.2)$ $(0.05,0.3)$
$l>L$ $(0.05,0.2)$ $(0.05,0.2)$
target boundary $T_{l}$ and Huber threshold $\delta_{l}$, we show the configuration in Table˜2 (L is set to 30 in Emu3 and 29 in Janus-Pro). We assign distinct values of T to different layers of the target model while dynamically adjusting $\delta$ based on the layer-wise similarity between Emu3, Janus-Pro and target models. This adaptive strategy prevents performance degradation caused by excessively rigid constraints. The AIA loss will combine with next-token-prediction loss, term $\mathcal{L}_{\text{NTP}}$ as:
where $\lambda$ is set to 40 in the experiment by default.
3.3 Training Details Under Different Architectures
Emu3. We load the pretrained (PT) weights of Emu3 and perform SFT training using our own data. Note that Emu3 only employs unified training at this stage; the results reported in the original paper were obtained through separate training. Since Emu3’s performance at the pretraining stage is relatively poor, as shown in fig.˜3, we incorporate the AIA loss with varying weights during the SFT stage and observe that the NTP loss convergence trends remain nearly identical. This indicates that, when the learning target is set appropriately, incorporating the AIA loss during SFT does not significantly affect the model’s pretrained knowledge.
Janus-Pro. Since Janus-Pro only provides the final SFT weights, we perform post-training on this basis. When data quality does not differ significantly, this greatly increases the difficulty of tuning, as the model’s distribution is already highly fixed at this stage (i.e., a point further supported by the variance comparison between Emu3 and Janus-Pro in Table˜1). However, this also aligns more closely with realistic settings: how to perform fine-grained tuning when only the final weights are accessible and attention distribution adjustment is required. As shown in fig.˜3, the model is highly sensitive to the weight of the AIA loss at this stage, yet it can still achieve the desired effect when $\lambda$ is set appropriately. We further validate this in section˜4.3.
@@ 3.3 Training Details Under Different Architectures
Gen. Only
SDXL Podell et al. (2023) - Diff - 0.55 74.65
SD3-medium Esser et al. (2024) 2B Diff - 0.74 84.08
Infinity Han et al. (2025) 2B VAR - 0.73 83.50
Infinity Han et al. (2025) 8B VAR - 0.79 86.60
FLUX.1-dev Labs (2024) 12B Diff - 0.82 84.00
Emu3-Gen Wang et al. (2024) 8B AR - 0.66 80.60
Qwen-Image Wu et al. (2025a) 7B+20B AR+Diff - 0.87 88.32
Und. Only
Emu3-Chat Wang et al. (2024) 8B AR 31.6 58.5 36.6 37.2 85.2 1244 - -
Qwen2.5-VL Bai et al. (2025b) 3B AR 53.1 79.1 - 61.8 - - - -
Qwen2.5-VL Bai et al. (2025b) 7B AR 58.6 83.5 - 66.6 - 1685 - -
InternVL2.5 Chen et al. (2024) 8B AR 56.2 84.6 - 62.8 90.6 - - -
InternVL3 Zhu et al. (2025) 8B AR 62.7 83.4 - 81.3 91.1 - - -
Qwen3-VL Bai et al. (2025a) 8B AR 69.6 85.0 - - - - - -
Uni. Frozen MLLM
MetaQuery-XL Pan et al. (2025) 7B+1.6B AR+Diff 58.6 83.5 - 66.6 - 1685 0.80 82.05
Blip3-o Chen et al. (2025a) 7B+1.4B AR+Diff 58.6 83.5 - 66.6 - 1685 0.84 81.60
UniWorld-V1 Lin et al. (2025) 7B+12B AR+Diff 58.6 83.5 - 67.1 - 1685 0.84 81.38
OmniGen2 Wu et al. (2025b) 3B+4B AR+Diff 53.1 79.1 - 61.8 - - 0.86 83.57
Uni. MoE/MoT Arch
BAGEL Deng et al. (2025) 7B+7B AR+Diff 55.3 85.0 69.3 67.2 - 1687 0.88 85.07
OneCat Li et al. (2025) 3B+3B+3B AR+VAR 41.9 78.8 - 52.2 - 1630 0.90 84.53
Uni. Double Image Encoders / Training Objectives
Show-o Xie et al. (2025b) 1.3B AR+Mask 26.7 - - - 80.0 1097 0.69 67.27
Show-o2 Xie et al. (2025c) 7B AR+Diff 48.9 79.3 - - - 1621 0.76 86.14
Janus-Pro Chen et al. (2025c) 7B AR 41.0 65.54 (79.2) 47.3 50.0 87.4 1567 0.80 84.19
@@
Uni. Purely
Chameleon Team (2024) 7B AR 28.4 35.7 - 8.3 - - 0.39 -
VILA-U Wu et al. (2024c) 7B AR 32.2 66.6 22.0 27.7 83.9 1336 0.39 72.48
Emu3 Wang et al. (2024) $\dagger$ 8B AR 31.6 61.4 8.7 15.1 77.3 910 0.60 79.24
@@
w/o stage-level intensity 40.2 67.4 43.1 44.0 87.6 1543.2 0.79 83.79
w/o Huber 41.2 73.1 47.6 47.3 88.2 1613.9 0.80 84.25
w/o AIA (baseline) 40.7 71.5 47.3 49.2 88.1 1593.1 0.80 84.19
AIA vs. Knowledge Distillation (Janus-Pro)
AIA (ours) 42.1 75.6 48.0 49.8 89.8 1656.4 0.81 84.49
Feature-level 32.5 58.6 41.0 40.3 85.5 1485.4 0.72 81.97
Output-level 40.2 68.4 47.3 46.9 87.7 1603.7 0.79 83.97
@@
Qwen3-VL+FLUX 40.2 72.1 45.5 46.0 87.5 1555.1 0.76 83.56
Qwen3-VL+SimpleAR 40.3 72.5 44.1 46.6 88.0 1546.7 0.75 82.89
@@
Qwen3-VL+Qwen-Image 41.5 74.3 47.3 48.6 89.5 1623.4 0.80 84.12
@@
Gen:Und: 2:1 41.8 74.2 46.6 49.3 89.1 1643.6 0.81 84.23
Gen:Und: 4:1 41.2 72.8 47.3 48.5 88.3 1621.4 0.80 83.87
Gen:Und: 1:2 42.0 74.8 47.3 49.3 89.3 1645.9 0.79 83.12